OpenStudy (goformit100):

Find the coordinates of the point which divides the line segment joining the points (1, –2, 3) and (3, 4, –5) in the ratio 2 : 3 (i) internally, and (ii) externall

OpenStudy (goformit100):

@terenzreignz

terenzreignz (terenzreignz):

What do internally and externally mean?

OpenStudy (agent0smith):

Probably able to use vectors again... find the vector from one point to the other, then multiply it by 2/5 or 3/5 (guessing that's internally/externally), then find it's endpoint.

OpenStudy (aravindg):

if internally use +sign ..externally -ve sign

OpenStudy (agent0smith):

If the points are A and B, find the endpoint by adding 2/5 of the vector AB, to point A

terenzreignz (terenzreignz):

Possibly similar to the midpoint formula, which goes \[\huge \left(\frac12(x_1+x_2),\frac12(y_1+y_2)\frac12(z_1+z_2)\right)\] Except instead of \(\large \frac12\), you multiply \(\large \frac25\) instead.

OpenStudy (agent0smith):

^that looks like it should work too (looks kinda like the result of what i said, just from trying to do it in my head)

OpenStudy (goformit100):

OH so we can also solve it using determinant or what ?

terenzreignz (terenzreignz):

Determinants? This looks too simple to have to resort to determinants... :/

OpenStudy (goformit100):

oh

terenzreignz (terenzreignz):

On second thought, that formula above may be too good to be true...

OpenStudy (agent0smith):

With vectors I think it'd be \[\large (x _{1}+\frac{ 2 }{ 5 }(x _{2}-x _{1}), y _{1}+\frac{ 2 }{ 5 }(y _{2}-y _{1}), z _{1}+\frac{ 2 }{ 5 }(z _{2}-z _{1}))\] Hopefully that's right, I'm too lazy to check on paper. It doesn't look exactly like @terenzreignz formula...

terenzreignz (terenzreignz):

Yeah... that's more like... It's just that when you take \(\frac12\) it should result into the more nice-looking midpoint formula. Yikes... I'm losing my touch :D

OpenStudy (agent0smith):

Yeah I assumed yours would work too... but it only works with half because 1-1/2 = 1/2

terenzreignz (terenzreignz):

Yup :) I learned.

OpenStudy (agent0smith):

Me too :) i figured that might be a handy shortcut for finding points not halfway between two points.

OpenStudy (agent0smith):

I guess externally, like @AravindG said, would just be subtracting 2/5. It doesn't seem to specify from which point to start from, though, and there's multiple possible points depending on which point you take the 2:5 ratio from.

OpenStudy (agent0smith):

@terenzreignz and @goformit100 there is actually a formula for this question (it's essentially what i posted above): http://www.learnnext.com/lesson/CBSE-X-Maths-Section-Formula.htm

OpenStudy (agent0smith):

It only gives for 2 dimensions but you just need to add the third in the same way.